The sum of three numbers is 23. The second number is one more than three times the first,and the sum of the first and third is 16. Find the numbers.
Answer by ewatrrr(24785)


Hi,
Let x, (3x+1), the second, and (16-x), the third, represent the numbers.
Question states thier sum is 23***
x + (3x+1) + (16 -x) = 23
solving for x
3x + 17 = 23
3x = 6
x = 2
the numbers are 2, 7 (3*2+1) and 14 (16-2)
CHECKING our Anwser***
2 + 7 + 14 = 23